# What is the Relationship between Tensor Factorizations and Circuits (and How Can We Exploit it)?

### v1

@@ 8 Conclusion
We hope this can propel research on how to design more and more scalable low-rank parameterization for probabilistic inference. To this end, we highlighted a number of possible future venues for the matrix and tensor factorization communities that leverage the connection with circuits we established: designing hierarchical factorizations with non-tree structures (Section 2.2); using the property-driven calculus that circuits offer to automatically derive tractable algorithms in a compositional way (Section 2.2); treat non-negative (hierarchical) factorizations as deep latent variable models (Section 3.1); devise factorizations over non-discrete and non-linear input spaces (Section 3.2); embed logical constraints to realize neuro-symbolic systems that can reason with symbolic knowledge (Section 3.3); devising alternative ways to compactly encode distributions, going beyond probability masses or densities (Section 3.4); as well as devising flexible factorizations by changing only the structure of (some) layers in a circuit representation (Section 5.2).
From the point of view of the circuit community, we leveraged this connection to systematize and demystify the construction of modern tensorized and overparameterized circuits (Section 4). We proposed a single pipeline that generalizes existing (tensor factorization and circuit) architectures and introduced a new nomenclature, based on the steps of our pipeline, to understand old but also new architectures that can be created by mixing & matching these steps (see Table 1). Our empirical analysis of popular ways to combine these ingredients highlights how lower-rank structures can be easier to learn and useful to compress higher-rank layers (Section 7). Finally, we distilled our findings in clear-cut recommendations (Sections 7, 7 and 7) for practitioners that want to learn and scale circuits on high-dimensional data, and we hope this can foster future rigorous analysis.
@@
This work is fundamental research in probabilistic modeling and reasoning and as such the algorithms and architectures discussed here can impact many possible downstream applications, in ways that go beyond our control.
For example, circuits, or tensor factorizations, could be used in computer vision classifiers to amplify the bias already encoded in non-curated datasets or be used in safety-critical applications without eliciting valid safety requirements. Since it is hard to foresee all possible future misuses, we urge practitioners to pay attention to concrete problematic uses of our methodologies: use the time that tractable models save while performing inference to reflect on the direct impact your application can have.
@@
AV and AM devised the original idea of tensorizing and compressing circuits in a unified pipeline. This has yielded a preliminary workshop paper whose content can be traced to Sections 4 and 5. AV and GV wrote the workshop paper and AM designed and run its experiments, which constituted the basis for the experiments in Section 8. LL traced the theoretical connections with tensor factorizations and tensor networks, and lead the writing of all sections in the journal paper with AV. AV and LL are responsible for designing and drawing all circuit pictures in TikZ, with the exception of Fig. 15 done by GG who also provided feedback for several plots. GG simplified the algorithmic pipeline, wrote the sampling algorithm and designed new experiments. GG and AM re-run and extended the workshop experiments to adapt them to the journal version, and GG wrote Section 7 with AV. EQ, RP and CdC critically read the manuscript. AV supervised all the phases of the project and provided feedback.
@@ Acknowledgments
We would like to acknowledge Nicolas Gillis and Grigorios Chrysos for providing feedback on an early draft of the paper. AV wants to thank Iain Murray for forwarding an email from Raul Garcia-Patron Sanchez inquiring about tensor networks and ML, and the latter for many following discussions about tensor networks. The Eindhoven University of Technology authors thank the support from the Eindhoven Artificial Intelligence Systems Institute and the Department of Mathematics and Computer Science of TU Eindhoven. This project has received funding from the European Union’s EIC Pathfinder Challenges 2022 programme under grant agreement No 101115317 (NEO). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Innovation Council. Neither the European Union nor the European Innovation Council can be held responsible for them. AV was supported by the “UNREAL: a Unified Reasoning Layer for Trustworthy ML” project (EP/Y023838/1) selected by the ERC and funded by UKRI EPSRC.
@@ A.1 Tucker as a Circuit
Let $\bm{\mathcal{T}}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ be a tensor being decomposed via a multilinear rank-$(R_{1},\ldots,R_{d})$ Tucker factorization, as in Eq. 2. Then, there exists a circuit $c$ over variables $\bm{\mathrm{X}}=\{X_{j}\}_{j=1}^{d}$ with $\mathsf{dom}(X_{j})=[I_{j}]$, $j\in[d]$ computing the same factorization. Moreover, we have that $|c|\in\mathcal{O}(\prod_{j=1}^{d}R_{j})$.
@@
Note that in Eq. 21 each product $c_{1,r_{1}}^{\mathsf{in}}(x_{1})\cdots c_{d,r_{d}}^{\mathsf{in}}(x_{d})$ can be computed by a product unit $c_{r_{1}\cdots r_{d}}^{\mathsf{prod}}$ over variables $\bm{\mathrm{X}}$. Moreover, we encode each $c_{j,r_{j}}^{\mathsf{in}}$ as an input unit, for all $j\in[d]$ and $r_{j}\in[R_{j}]$. In addition, the collection of sums $\sum_{r_{1}=1}^{R_{1}}\cdots\sum_{r_{d}=1}^{R_{d}}$ that are weighted by the $w_{r_{1}\cdots r_{d}}$ can be computed by a single sum unit having $\prod_{j=1}^{d}R_{j}$ inputs, i.e., the products $c_{r_{1}\cdots r_{d}}^{\mathsf{prod}}$ with $r_{j}\in[R_{j}]$ for all $j$. Therefore, we have that the overall circuit size is $|c|\in\mathcal{O}(\prod_{j=1}^{d}R_{j})$. Finally, we take $w_{r_{1}\cdots r_{d}}$ as the entries of the core tensor $\bm{\mathcal{W}}$, and make each input unit $c_{j,r_{j}}^{\mathsf{in}}$ compute $c_{j,r_{j}}^{\mathsf{in}}(x_{j})=v_{x_{j},r_{j}}^{(j)}$ for the factor matrices $\{\bm{\mathrm{V}}^{(j)}\}_{j=1}^{d}$ in the Tucker factorization. That is, $c(\bm{\mathrm{X}})$ computes the same Tucker factorization given by hypothesis. ∎
@@ Appendix B Many Tensorized PC Architectures can be Obtained through our Pipeline
Born machines (BM) (Han et al., 2018) and Tensor-Train Density Estimators (TTDE) (Novikov et al., 2021) are probabilistic models used to estimate probability mass functions and probability density functions, respectively. They are obtained by efficiently squaring an MPS, which is a structured decomposable tensorized circuit as for 3. Note that such a tensorized circuit can be obtained using the same region graph and tensorized layer used to construct a non-negative MPS, but instead just relax the non-negativity assumption over its parameters. It is known that squaring a MPS (resp. a structured decomposable tensorized circuit) yields a BM (resp. another structured decomposable tensorized circuit having the same layers but with a quadratic width increase). See e.g. Proposition 3 in Loconte et al. (2024a). Therefore, BMs and TTDEs can be retrieved through our circuit construction pipeline by overaparameterizing a linear-tree region graph (LT) with $\text{CP}^{\top}$layers, followed by efficiently squaring the resulting circuit (Vergari et al., 2021).
Squared non-monotonic PCs ($\text{NPC}^{2}$) (Loconte et al., 2024a) are generalizations of BMs and TTDEs that also include the squaring of tensorized circuits obtained by overparameterizing a random binary tree region graph (as in RAT-SPNs above), as well as using Tucker layers instead of $\text{CP}^{\top}$layers. Furthermore, the original implementation of $\text{NPC}^{2}$ allows circuits to be folded.

### v2

@@ 8 Conclusion
We hope this can propel research on how to design more and more scalable low-rank parameterization for probabilistic inference. To this end, we highlighted a number of possible future avenues for the matrix and tensor factorization communities that leverage the connection with circuits we established: designing hierarchical factorizations with non-tree structures (Section 2.2); using the property-driven calculus that circuits offer to automatically derive tractable algorithms in a compositional way (Section 2.2); treat non-negative (hierarchical) factorizations as deep latent variable models (Section 3.1); devise factorizations over non-discrete and non-linear input spaces (Section 3.2); embed logical constraints to realize neuro-symbolic systems that can reason with symbolic knowledge (Section 3.3); devising alternative ways to compactly encode distributions, going beyond probability masses or densities (Section 3.4); as well as devising flexible factorizations by changing only the structure of (some) layers in a circuit representation (Section 5.2).
From the point of view of the circuit community, we leveraged this connection to systematize and demystify the construction of modern tensorized and overparameterized circuits (Section 4). We proposed a single pipeline that generalizes existing (tensor factorization and circuit) architectures and introduced a new nomenclature, based on the steps of our pipeline, to understand old but also new architectures that can be created by mixing & matching these steps (see Table 1). Our empirical analysis of popular ways to combine these ingredients highlights how lower-rank structures can be easier to learn and useful to compress higher-rank layers (Section 6). Finally, we distilled our findings in clear-cut recommendations (Sections 6, 6 and 6) for practitioners that want to learn and scale circuits on high-dimensional data, and we hope this can foster future rigorous analysis.
@@
This work is fundamental research in probabilistic modeling and reasoning and as such the algorithms and architectures discussed here can impact many possible downstream applications, in ways that go beyond our control. For example, circuits, or tensor factorizations, could be used in computer vision classifiers to amplify the bias already encoded in non-curated datasets or be used in safety-critical applications without eliciting valid safety requirements. Since it is hard to foresee all possible future misuses, we urge practitioners to pay attention to concrete problematic uses of our methodologies: use the time that tractable models save while performing inference to reflect on the direct impact your application can have.
@@
AV and AM devised the original idea of tensorizing and compressing circuits in a unified pipeline. This has yielded a preliminary workshop paper whose content can be traced to Sections 4 and 5. AV and GV wrote the workshop paper and AM designed and run its experiments, which constituted the basis for the experiments in Section 6. LL traced the theoretical connections with tensor factorizations and tensor networks, and lead the writing of all sections in the journal paper with AV. AV and LL are responsible for designing and drawing all circuit pictures in TikZ, with the exception of Fig. 16 done by GG who also provided feedback for several plots. GG simplified the algorithmic pipeline, wrote the sampling algorithm in Appendix C, produced Appendix F, and designed new experiments. GG and AM re-run and extended the workshop experiments to adapt them to the journal version, and GG wrote Section 6 with AV. EQ, RP and CdC critically read the manuscript. AV supervised all the phases of the project and provided feedback.
@@ Acknowledgments
We would like to acknowledge Nicolas Gillis and Grigorios Chrysos for providing feedback on an early draft of the paper. AV wants to thank Iain Murray for forwarding an email from Raul Garcia-Patron Sanchez inquiring about tensor networks and ML, and the latter for many following discussions about tensor networks. The Eindhoven University of Technology authors thank the support from (i) the Eindhoven Artificial Intelligence Systems Institute, (ii) the Department of Mathematics and Computer Science of TU Eindhoven, and (iii) the EU European Defence Fund Project KOIOS (EDF-2021-DIGIT-R-FL-KOIOS). This project has received funding from the European Union’s EIC Pathfinder Challenges 2022 programme under grant agreement No 101115317 (NEO). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Innovation Council. Neither the European Union nor the European Innovation Council can be held responsible for them. AV was supported by the “UNREAL: a Unified Reasoning Layer for Trustworthy ML” project (EP/Y023838/1) selected by the ERC and funded by UKRI EPSRC.
@@ A.1 Tucker as a Circuit
Let $\bm{\mathcal{T}}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ be a tensor being decomposed via a multilinear rank-$(R_{1},\ldots,R_{d})$ Tucker factorization, as in Eq. 2. Then, there exists a circuit $c$ over variables $\bm{\mathrm{X}}=\{X_{j}\}_{j=1}^{d}$ with $\mathsf{dom}(X_{j})=[I_{j}]$, $j\in[d]$ computing the same factorization. Moreover, we have that $|c|\in\mathcal{O}(d\prod_{j=1}^{d}R_{j})$.
@@
Note that in Eq. 19 each product $c_{1,r_{1}}^{\mathsf{in}}(x_{1})\cdots c_{d,r_{d}}^{\mathsf{in}}(x_{d})$ can be computed by a product unit $c_{r_{1}\cdots r_{d}}^{\mathsf{prod}}$ over variables $\bm{\mathrm{X}}$. Moreover, we encode each $c_{j,r_{j}}^{\mathsf{in}}$ as an input unit, for all $j\in[d]$ and $r_{j}\in[R_{j}]$. In addition, the collection of sums $\sum_{r_{1}=1}^{R_{1}}\cdots\sum_{r_{d}=1}^{R_{d}}$ that are weighted by the $w_{r_{1}\cdots r_{d}}$ can be computed by a single sum unit having $\prod_{j=1}^{d}R_{j}$ inputs, i.e., the products $c_{r_{1}\cdots r_{d}}^{\mathsf{prod}}$ with $r_{j}\in[R_{j}]$ for all $j$. Since each product units has $d$ inputs, we have that the overall circuit size is $|c|\in\mathcal{O}(d\prod_{j=1}^{d}R_{j})$. Finally, we take $w_{r_{1}\cdots r_{d}}$ as the entries of the core tensor $\bm{\mathcal{W}}$, and make each input unit $c_{j,r_{j}}^{\mathsf{in}}$ compute $c_{j,r_{j}}^{\mathsf{in}}(x_{j})=v_{x_{j},r_{j}}^{(j)}$ for the factor matrices $\{\bm{\mathrm{V}}^{(j)}\}_{j=1}^{d}$ in the Tucker factorization. That is, $c(\bm{\mathrm{X}})$ computes the same Tucker factorization given by hypothesis. ∎
@@ Appendix B Many Tensorized PC Architectures can be Obtained through our Pipeline
Born machines (BM) (Han et al., 2018) and Tensor-Train Density Estimators (TTDE) (Novikov et al., 2021) are probabilistic models used to estimate probability mass functions and probability density functions, respectively. They are obtained by efficiently squaring an MPS, which is a structured decomposable tensorized circuit as for 3. Note that such a tensorized circuit can be obtained using the same region graph and tensorized layer used to construct a non-negative MPS, but instead just relax the non-negativity assumption over its parameters. It is known that squaring a MPS (resp. a structured decomposable tensorized circuit) yields a BM (resp. another structured decomposable tensorized circuit having the same layers but with a quadratic width increase). See e.g. Proposition 3 in Loconte et al. (2024). Therefore, BMs and TTDEs can be retrieved through our circuit construction pipeline by overaparameterizing a linear-tree region graph (LT) with $\text{CP}^{\top}$layers, followed by efficiently squaring the resulting circuit (Vergari et al., 2021).
Squared non-monotonic PCs ($\text{NPC}^{2}$) (Loconte et al., 2024) are generalizations of BMs and TTDEs that also include the squaring of tensorized circuits obtained by overparameterizing a random binary tree region graph (as in RAT-SPNs above), as well as using Tucker layers instead of $\text{CP}^{\top}$layers. Furthermore, the original implementation of $\text{NPC}^{2}$ allows circuits to be folded.
@@ Appendix E Additional Results
E.1 Results on UCI Tabular Datasets
Number of samples
$D$ train validation test
Power $6$ $1{,}659{,}917$ $184{,}435$ $204{,}928$
Gas $8$ $852{,}174$ $94{,}685$ $105{,}206$
Hepmass $21$ $315{,}123$ $35{,}013$ $174{,}987$
MiniBooNE $43$ $29{,}556$ $3{,}284$ $3{,}648$
BSDS300 $63$ $1{,}000{,}000$ $50{,}000$ $250{,}000$
Table E.5: UCI dataset statistics. Dimensionality $D$ and number of samples of each dataset split after the preprocessing by Papamakarios et al. (2017).
Density estimation on tabular datasets.
Following Papamakarios et al. (2017), we evaluate our tensorized architectures for density estimation on five tabular datasets. For each dataset, we randomly construct 8 binary tree region graphs (cf. Section 4.1), and build a mixture of tensorized PCs based of them. Specifically, following our mix-and-match approach Table 1, we build RND-CP and RND-Tucker architectures which we run for several model sizes $K$ and learning rates (see below). Differently from images, all these datasets contain continuous features, which we model using input layers encoding Gaussian likelihoods. We train all PCs for up to 1000 epochs or until convergence, using Adam as optimizer and 512 as batch size. Furthermore, we perform the experiments using three different learning rates: $10^{-3}$, $5\cdot 10^{-3}$, and $10^{-2}$, and report the best results according to the validation set log-likelihood.
Results.
Fig. E.5 reports the best results from our models, where we see that Tucker layers outperform CP layers on the two lowest dimensional datasets – Power and Gas – which also have the highest number of training data points (see Table E.5). On the other hand, CP-based architectures outperform Tucker-based ones on the other three datasets (Hepmass, MiniBooNE and BSDS300), even though the latter have a much higher number of trainable parameters then the former for a fixed $K$ (i.e., $K^{2}$ for CP while $K^{3}$ for Tucker). Our results suggest that the more aggressive over-parameterization of Tucker layers lead to a more difficult optimization for high-dimensional datasets and thus for deeper tensorized PCs.
Power Gas Hepmass M.BooNE BSDS300
MADE -3.08 3.56 -20.98 -15.59 148.85
RealNVP 0.17 8.33 -18.71 -13.84 153.28
MAF 0.24 10.08 -17.73 -12.24 154.93
NSF 0.66 13.09 -14.01 -9.22 157.31
Gaussian -7.74 -3.58 -27.93 -37.24 96.67
EiNet-LRS 0.36 4.79 -22.46 -34.21
TTDE 0.46 8.93 -21.34 -28.77 143.30
RND-CP 0.28 5.01 -22.52 -30.69 120.82
RND-Tucker 0.52 8.41 -23.47 -31.30 119.09
Figure E.5: Tucker layers are harder to scale than CP layers on high-dimensional UCI datasets. The right table is the one reported in Table 3. The left plots show the train and test log-likelihoods of our architectures as the size $K$ of the layers increases. We observe that increasing $K$ is generally beneficial for CP layers in all UCI datasets (left). However, increasing $K$ in Tucker layers can decrease performances for higher-dimensional datasets, as shown for the cases of Hepmass and MiniBooNE. The left plots showing the train set log-likelihoods (dotted lines) are evidence that the decrease of performances of tensorized PCs with Tucker layers is not due to overfitting.
Appendix F How to implement (folded) layers?
In this section, we provide pytorch snippets to implement a folded Categorical input layer, as well as all the folded sum-product layers we introduced, i.e.  Tucker (Eq. Tucker-layer), CP (Eq. CP-layer), $\text{CP}^{\top}$ (Eq. $\text{CP}^{\top}$-layer), CPS and CPXS (Eq. CPS-layer). Our tensorized circuit architectures are nothing more than a sequential application of such layers. We will release our code upon acceptance.
1def categorical_layer(batch, param):
2 """EvaluatesafoldedCategoricallayerwithCcategories.
3
4␣␣␣␣:parambatch:Inputbatch(num_folds,num_channels,batch_size)
5␣␣␣␣:paramparam:Rawparameters(num_folds,K,num_channels,C)
6␣␣␣␣:returnlog_prob:Outputlogprob(num_folds,K,batch_size)
7␣␣␣␣"""
8 cat_log_prob = param.log_softmax(-1)
9 idx_fold = torch.arange(batch.size(0))[:, None, None]
10 idx_channel = torch.arange(batch.size(1))[None, :, None]
11 log_prob = cat_log_prob[idx_fold, :, idx_channel, batch].sum(dim=1)
12 return log_prob.transpose(-1, -2)
Figure F.1: Pytorch snippet for a folded categorical layer. The snippet details the evaluation of a folded categorical layer, which can be used for any type of categorical variable. For instance, it can be used to evaluate the log-likelihood of pixels in image modeling as well as tokens in language modeling. Note how the raw input parameters (param) undergo a log-softmax reparameterization (Section 3.2) so as to model valid log-probabilities for many categorical distributions.
1def tucker_layer(log_prob, W):
2 """EvaluatesafoldedTuckerlayerwitharity2vialog-sum-exp.
3
4␣␣␣␣:paramlog_prob:Inputlogprob(num_folds,2,K,batch_size)
5␣␣␣␣:paramW:Layerweightsinprobdomain(num_folds,K,K,O)
6␣␣␣␣:returnout_log_prob:Outputlogprob(num_folds,O,batch_size)
7␣␣␣␣"""
8 max_log_prob = log_prob.max(dim=2, keepdim=True).values
9 exp_log_prob = torch.exp(log_prob - max_log_prob)
10 out_log_prob = max_log_prob.sum(dim=1) + torch.einsum(
11 "fib,fjb,fijo->fob",
12 exp_log_prob[:, 0], exp_log_prob[:, 1], W).log()
13 return out_log_prob
Figure F.2: Pytorch snippet for a folded Tucker layer. A folded Tucker layer $\bm{\mathrm{\ell}}$ is parameterized by a tensor $\bm{\mathcal{W}}$ of shape $F\times O\times K^{2}$, and computes $F$ Tucker layer (Eq. Tucker-layer) $\{\bm{\mathrm{\ell}}^{(n)}\}_{n=1}^{F}$ in parallel. Specifically, the layer $\bm{\mathrm{\ell}}$ computes
                              $\bm{\mathrm{\ell}}\left(\bigcup\nolimits_{n=1}^{F}\bm{\mathrm{Y}}^{(n)}\right)_{n:}=\bm{\mathrm{W}}_{n::}\left[\bm{\mathrm{\ell}}_{1}\left(\bigcup\nolimits_{n=1}^{F}\bm{\mathrm{Z}}_{1}^{(n)}\right)_{n:}\otimes\bm{\mathrm{\ell}}_{2}\left(\bigcup\nolimits_{n=1}^{F}\bm{\mathrm{Z}}_{2}^{(n)}\right)_{n:}\right],$
where $\bm{\mathrm{\ell}}_{1}$ (resp. $\bm{\mathrm{\ell}}_{2}$) denotes a folded layer computing the $F$ left (resp. right) inputs to $\bm{\mathrm{\ell}}^{(n)}$, each defined over variables $\bm{\mathrm{Z}}_{1}^{(n)}$ (resp. $\bm{\mathrm{Z}}_{2}^{(n)}$), and $\bm{\mathrm{W}}_{n::}\in\mathbb{R}^{O\times K^{2}}$ is the parameter matrix of $\bm{\mathrm{\ell}}^{(n)}$. Note that the snippet shapes the tensor $\bm{\mathcal{W}}$ as $F\times K\times K\times O$ for convenience with the einsum operation.
1def cp_layer(log_prob, W1, W2):
2 """EvaluatesafoldedCPlayerwitharity2vialog-sum-exp.
3
4␣␣␣␣:paramlog_prob:Inputlogprobs,(num_folds,2,K,batch_size)
5␣␣␣␣:paramW1:Foldedlayerweights,(num_folds,K,O)
6␣␣␣␣:paramW2:Foldedlayerweights,(num_folds,K,O)
7␣␣␣␣:returnout_log_prob:Outputlogprobs,(num_folds,O,batch_size)
8␣␣␣␣"""
9 max_log_prob = log_prob.max(dim=2, keepdim=True).values
10 exp_log_prob = torch.exp(log_prob - max_log_prob)
11 out_log_prob = max_log_prob.sum(dim=1) + torch.einsum(
12 "fib,fjb,fio,fjo->fob",
13 exp_log_prob[:, 0], exp_log_prob[:, 1], W1, W2).log()
14 return out_log_prob
Figure F.3: Pytorch snippet for a folded CP layer. A folded CP layer $\bm{\mathrm{\ell}}$ is parameterized by two equally-sized tensors $\bm{\mathcal{W}}^{(1)}$ and $\bm{\mathcal{W}}^{(2)}$ of shape $F\times O\times K$, and computes $F$ CP layer (Eq. CP-layer) $\{\bm{\mathrm{\ell}}^{(n)}\}_{n=1}^{F}$ in parallel. Specifically, the layer $\bm{\mathrm{\ell}}$ computes
                           $\bm{\mathrm{\ell}}\left(\bigcup\nolimits_{n=1}^{F}\bm{\mathrm{Y}}^{(n)}\right)_{n:}=\left(\bm{\mathrm{W}}_{n::}^{(1)}\,\bm{\mathrm{\ell}}_{1}\left(\bigcup\nolimits_{n=1}^{F}\bm{\mathrm{Z}}_{1}^{(n)}\right)_{n:}\right)\odot\left(\bm{\mathrm{W}}_{n::}^{(2)}\,\bm{\mathrm{\ell}}_{2}\left(\bigcup\nolimits_{n=1}^{F}\bm{\mathrm{Z}}_{2}^{(n)}\right)_{n:}\right),$
where $\bm{\mathrm{\ell}}_{1}$ (resp. $\bm{\mathrm{\ell}}_{2}$) denotes a folded layer computing the $F$ left (resp. right) inputs to $\bm{\mathrm{\ell}}^{(n)}$, each defined over variables $\bm{\mathrm{Z}}_{1}^{(n)}$ (resp. $\bm{\mathrm{Z}}_{2}^{(n)}$), and $\bm{\mathrm{W}}^{(1)}_{n::},\bm{\mathrm{W}}^{(2)}_{n::}\in\mathbb{R}^{O\times K}$ are the parameter matrices of $\bm{\mathrm{\ell}}^{(n)}$.
1def cpt_layer(log_prob, W):
2 """EvaluatesafoldedCP-TlayerwitharityHvialog-sum-exp.
3
4␣␣␣␣:paramlog_prob:Inputlogprob(num_folds,H,K,batch_size)
5␣␣␣␣:paramW:Foldedlayerweights(num_folds,K,O)
6␣␣␣␣:returnout_log_prob:Outputlogprob(num_folds,O,batch_size)
7␣␣␣␣"""
8 log_prob = log_prob.sum(dim=1, keepdim=True)
9 max_log_prob = log_prob.max(dim=2, keepdim=True).values
10 exp_log_prob = torch.exp(log_prob - max_log_prob)
11 out_log_prob = max_log_prob.sum(dim=1) + torch.einsum(
12 "fib,fio->fob",
13 exp_log_prob[:, 0], W).log()
14 return out_log_prob
Figure F.4: Pytorch snippet for a folded $\text{CP}^{\top}$layer. A folded $\text{CP}^{\top}$layer $\bm{\mathrm{\ell}}$ is parameterized by tensor $\bm{\mathcal{W}}$ of shape $F\times O\times K$, and computes $F$ $\text{CP}^{\top}$layer (Eq. $\text{CP}^{\top}$-layer) $\{\bm{\mathrm{\ell}}^{(n)}\}_{n=1}^{F}$ in parallel. Specifically, the layer $\bm{\mathrm{\ell}}$ computes
                           $\bm{\mathrm{\ell}}\left(\bigcup\nolimits_{n=1}^{F}\bm{\mathrm{Y}}^{(n)}\right)_{n:}=\bm{\mathrm{W}}_{n::}\left(\,\bm{\mathrm{\ell}}_{1}\left(\bigcup\nolimits_{n=1}^{F}\bm{\mathrm{Z}}_{1}^{(n)}\right)_{n:}\odot\,\bm{\mathrm{\ell}}_{2}\left(\bigcup\nolimits_{n=1}^{F}\bm{\mathrm{Z}}_{2}^{(n)}\right)_{n:}\right),$
where $\bm{\mathrm{\ell}}_{1}$ (resp. $\bm{\mathrm{\ell}}_{2}$) denotes a folded layer computing the $F$ left (resp. right) inputs to $\bm{\mathrm{\ell}}^{(n)}$, each defined over variables $\bm{\mathrm{Z}}_{1}^{(n)}$ (resp. $\bm{\mathrm{Z}}_{2}^{(n)}$), and $\bm{\mathrm{W}}_{n::}\in\mathbb{R}^{O\times K}$ are the parameter matrices of $\bm{\mathrm{\ell}}^{(n)}$.
1def cpxs_layer(log_prob, W1, W2, D=None):
2 """EvaluatesaCP-XSlayerwitharity2vialog-sum-exp.
3␣␣␣␣␣␣␣IfDisNonethenevaluatesaCP-Slayer.
4
5␣␣␣␣:paramlog_prob:Inputlogprobs,(num_folds,2,K,batch_size)
6␣␣␣␣:paramW1:Foldedlayerweights,(K,O)
7␣␣␣␣:paramW2:Foldedlayerweights,(K,O)
8␣␣␣␣:paramD:Optionalweights(num_folds,O)
9␣␣␣␣:returnout_log_prob:Outputlogprobs,(num_folds,O,batch_size)
10␣␣␣␣"""
11 max_log_prob = log_prob.max(dim=2, keepdim=True).values
12 exp_log_prob = torch.exp(log_prob - max_log_prob)
13 out_log_prob = max_log_prob.sum(dim=1) + torch.einsum(
14 "fib,fjb,io,jo->fob",
15 exp_log_prob[:, 0], exp_log_prob[:, 1], W1, W2).log()
16 if D:
17 out_log_prob += D.log().unsqueeze(-1)
18 return out_log_prob
Figure F.5: Pytorch snippet for a CPS layer. A CPS layer $\bm{\mathrm{\ell}}$ is parameterized by two equally-sized matrices $\bm{\mathrm{W}}^{(1)}$ and $\bm{\mathrm{W}}^{(2)}$ of shape $O\times K$, and a matrix $\bm{\mathrm{D}}$ of shape $F\times O$. The layer computes $F$ CP layer (Eq. CPS-layer) $\{\bm{\mathrm{\ell}}^{(n)}\}_{n=1}^{F}$ in parallel, each parameterized by $\bm{\mathrm{W}}^{(1)}$ and $\bm{\mathrm{W}}^{(2)}$, and then weights their outputs by $\bm{\mathrm{D}}$. Specifically, the layer $\bm{\mathrm{\ell}}$ computes
                        $\bm{\mathrm{\ell}}\left(\bigcup\nolimits_{n=1}^{F}\bm{\mathrm{Y}}^{(n)}\right)_{n:}=\bm{\mathrm{d}}_{n:}\odot\left(\bm{\mathrm{W}}^{(1)}\,\bm{\mathrm{\ell}}_{1}\left(\bigcup\nolimits_{n=1}^{F}\bm{\mathrm{Z}}_{1}^{(n)}\right)_{n:}\right)\odot\left(\bm{\mathrm{W}}^{(2)}\,\bm{\mathrm{\ell}}_{2}\left(\bigcup\nolimits_{n=1}^{F}\bm{\mathrm{Z}}_{2}^{(n)}\right)_{n:}\right),$
where $\bm{\mathrm{\ell}}_{1}$ (resp. $\bm{\mathrm{\ell}}_{2}$) denotes a folded layer computing the $F$ left (resp. right) inputs to $\bm{\mathrm{\ell}}^{(n)}$, each defined over variables $\bm{\mathrm{Z}}_{1}^{(n)}$ (resp. $\bm{\mathrm{Z}}_{2}^{(n)}$), $\bm{\mathrm{W}}^{(1)},\bm{\mathrm{W}}^{(2)}\in\mathbb{R}^{O\times K}$ are the parameter matrices of $\bm{\mathrm{\ell}}^{(n)}$, and $\bm{\mathrm{D}}$ as a folded-dependent parameter matrix. When $\bm{\mathrm{D}}=\bm{1}$ we retrieve CPXS.